Find the unit tangent vector to any point on the curve
r=(t^2+1)I+(4t-3)j+(2t^2-6t)k
1
Expert's answer
2020-06-30T18:43:02-0400

Given "r =(1+t^2)\hat{i} + (4t-3)\hat{j} + (2t^2-6t)\hat{k}"

differentiating it with respect to t,

"r' = 2t\hat{i} + 4\hat{j} + (4t-6)\hat{k}"

unit tangent vector,

"\hat{r'} = \frac{\vec{r'}}{|r'|} = \frac{2t\hat{i} + 4\hat{j} + (4t-6)\hat{k}}{\sqrt{4t^2 + 16 + (4t-6)^2}}" "=\frac{2t\hat{i} + 4\hat{j} + (4t-6)\hat{k}}{\sqrt{(20t^2-12t+52)}} = \frac{t\hat{i} + 2\hat{j} + (2t-3)\hat{k}}{\sqrt{(5t^2-3t+13)}}"